TABLE 15-2 Metabolic Rates (65-kg human) For a time period of 24 hours; list all the food that you consume and the amount of calories that each food contains in Calories. Also monitor all physical activities for the same period of 24 hours. Include sleeping, sitting and account for 24 hours. Estimate the energy transformed using metabolic rates from Table 15-2, page 419 in your text. You can use other resources too. Compare and make conclusions. Metabolic Rate (approximate Activity kcal/h watts 60 70 115 200 230 Sleeping Sitting upright Light activity (eating, dressing. household chores) Moderate work (tennis, walking) Running (15 km/h) Bicycling (race) 100 12:00am-7:00am -sleep 7:00am-8:30-breakfast (25 cal) 8:30-10:00am -driving 10:00am- 3:00pm -sitting 3:00pm-4:30pm -driving 4:30pm-5:30pm- dinner (1,109 cal) 5:30-8:00pm- sitting 8:00pm-12:00am- sleep 400 460 1000 1100 1150 1270
Answer #1

1)5 hr of sleep= 5×60=300 kcal. lost

2)1.5 hr of eating=1.5×200=300 k cal lost

3)now sitting and driving will fall in sstting upright activity so from 8:30 to4:30 make it 8hr= 8×100=800 k cal lost

4)now 1.5 hr of dinner=1.5×200=300 k cal lost

5)now 2.5 hr of sitting =2.5×100=250 k cal lost

6) 4 hr of sleep =4×60=120 k cal

in total you have lost 2070 k cal

and taken only 25+1109=1134 cal